Two laser beams one of wave length 640 nm and the other 400 nm have same unit flux of photons. Their powers are in the ratio

Grade:12

1 Answers

Mayajyothi C J
7 years ago
P \alpha \frac{1}{\lambda ^{2}}
 
\frac{P_{1}}{P_{2}} =\frac{(\lambda _{2})^{2}}{(\lambda_{1})^{2}}
=\frac{400^{2}}{650^{2}}
=\frac{64}{169}